Lorena likes to play golf. The number of times per year that she plays depends on both the price of playing a round of golf as well as Lorenas income and the cost of other types of entertainment-in particular, how much it costs to go see a movie instead of playing golf. The three demand schedules in the table below show how many rounds of golf per year Lorena will demand at each price under three different scenarios. In scenario D, Lorenas income is $80,000 per year and movies cost $15 each. In scenario , Lorenas income is also $80,000 per year, but the price of seeing a movie rises to $17. And in scenario D3. Lorenas income goes up to $100,000 per year, while movies cost $17 Scenario D1 D2 D3 100,000 17 80,000 80,000 Income per year Price of movie ticket 15 S 17 S Price of Golf $55 $40 $25 15 25 40 Quantity Demanded 10 15 20 15 30 50
media%2Fc6c%2Fc6c418ce-1f65-4dc2-b10d-5f
media%2Ff76%2Ff76ae717-345a-4b4d-a8f0-83
Answer #1

a) Cross price Ed = % change in quantity demanded of golf / % change in price of movie tickets

Cross price Ed at $ 55 = [(10 - 15)/15 x 100]/[(17 - 15)/15 x 100] = - 33.3/13.33 = - 2.5

Cross price Ed at $ 40 = [(15 - 25)/25 x 100]/[(17 - 15)/15 x 100] = - 40/13.33 = - 3

Cross price Ed at $ 25 = [(20 - 40)/40 x 100]/[(17 - 15)/15 x 100] = - 50/13.33 = - 3.75

Cross price Ed is different at all prices.

Complementary goods.

b) Income Ed = % change in quantity demanded of golf / % change in the income

Income Ed at $ 55 = [(15 - 10)/10 x 100]/[(100,000 - 80,000)/80,000 x 100] = 50/25 = 2

Income Ed at $ 40 = [(30 - 15)/15 x 100]/[(100,000 - 80,000)/80,000 x 100] = 100/25 = 4

Income Ed at $ 25 = [(50 - 20)/20 x 100]/[(100,000 - 80,000)/80,000 x 100] = 150/25 = 6

Income Ed is different at all prices.

No, it is not inferior good, it is Normal good.